OpenStudy (anonymous):

assume that an angle B is an acute angle and tan(5B-9)= cot(7B-81degrees). find the angle B

OpenStudy (dumbcow):

\[5B-9 = \tan^{-1} (\frac{1}{\tan (7B-81)})\] \[5B-9 = \frac{1}{7B-81}\] \[35B^{2}-468B+728 = 0\] \[B \approx 1.8 , 11.6\]

OpenStudy (anonymous):

thank you so much!!!